Solve the following system of equations:
y=2x-5
y=-x+4
Your answer should be in the form of (x,y) coordinates.
Your answer should be in the format (x,y)

Answer by ikleyn(52781)

.
Solve the following system of equations:
y=2x-5
y=-x+4
Your answer should be in the form of (x,y) coordinates.
Your answer should be in the format (x,y)
~~~~~~~~~~~~~~~

Since left sides are identical, right sides should be equal

    2x-5 = -x+4.


Collect the terms with x on the left side; the constant terms on the right side

    2x + x = 5 + 4

      3x   =   9

       x   =   9/3 = 3.


Thus you just found value of x: it is 3.


To find y, substitute x= 3 into the second equation

    y = -x+4 = -3+4 = 1.


ANSWER.  The solution is  (x,y) = (3,1)  in the requested form.

Solved.